OpenStudy (anonymous):

Phillipa types at a rate of 50 wpm. A. How long would it take her to type a 500 word essay at this rate? B. How much longer would it take Kurt to type this essay if he types at 35 wpm?

OpenStudy (anonymous):

Easy , total words = rate * time use this equation and tell me ur answer

OpenStudy (anonymous):

isolate the unknown ( time) by taking rate to the other side :D

OpenStudy (anonymous):

50 * 500 = 2500?

OpenStudy (anonymous):

no , 50 is rate and 500 is total words , time = total words/rate

OpenStudy (anonymous):

But what about b?

OpenStudy (anonymous):

use same equation to find kurt's time , then find the difference

OpenStudy (anonymous):

Why difference?

OpenStudy (anonymous):

4 minutes 29 seconds

OpenStudy (anonymous):

29 thousandths of a second

OpenStudy (anonymous):

no u dont put it like that , kurt's time is 14 minutes and 10/35 , to convert to seconds , do 10/35 * 60

OpenStudy (anonymous):

but i subtracted their times

OpenStudy (anonymous):

yes but the seconds are wrong , 1min = 60seconds , according to what u did , 1min = 100 sec

OpenStudy (anonymous):

How do you convert? @JOELMATHEWS1234

OpenStudy (anonymous):

that's what i said , 500/35 = 14minutes and 10 / 35 minutes To convert 10/35 to seconds , multiply is by 60

OpenStudy (anonymous):

How did this ratio derive?

OpenStudy (anonymous):

500 / 35 , remainder is 10 , so put it in mixed fraction

OpenStudy (yttrium):

For A. Given: r = 50wpm total words = 500 Required: t \[Words _{total} = rt\] Therefore, \[t = Words _{total}/r\] t = 500words/50wpm t = 10 minutes

OpenStudy (yttrium):

For b Given: \[r _{2} = 35wpm\] total words = 500 Required: \[t _{2} - t _{1}\] Using the formula we used above, \[t _{2} = 500/35\] \[t _{2} = 14.28 minutes.\] \[t _{2} - t _{1} = 14.28 - 10 = 4.28 minutes\]